# Operator entanglement in $\mathrm{SU}(2)$-symmetric dissipative quantum many-body dynamics

### v2

@@ Abstract
The presence of symmetries can lead to nontrivial dynamics of operator entanglement in open quantum many-body systems, which characterizes the cost of an matrix product density operator (MPDO) representation of the density matrix in the tensor-network methods and provides a measure for the corresponding classical simulability. One example is the $\mathrm{U}(1)$-symmetric open quantum systems with dephasing, in which the operator entanglement increases logarithmically at late times instead of being suppressed by the dephasing. Here we numerically study the far-from-equilibrium dynamics of operator entanglement in a dissipative quantum many-body system with the more complicated $\mathrm{SU}(2)$ symmetry and dissipations beyond dephasing. We show that after the initial rise and fall, the operator entanglement also increases again in a logarithmic manner at late times in the $\mathrm{SU}(2)$-symmetric case. We find that this behavior can be fully understood from the corresponding $\mathrm{U}(1)$ subsymmetry by considering the symmetry-resolved operator entanglement. But unlike the $\mathrm{U}(1)$-symmetric case with dephasing, both the classical Shannon entropy associated with the probabilities for the half system being in different symmetry sectors and the corresponding symmetry-resolved operator entanglement have nontrivial contributions to the late time logarithmic growth of operator entanglement. Our results show evidence that the logarithmic growth of operator entanglement at long times is a generic behavior of dissipative quantum many-body dynamics with $\mathrm{U}(1)$ as the symmetry or subsymmetry and for more broad dissipations beyond dephasing. We show that the latter is valid even for open quantum systems with only $\mathrm{U}(1)$ symmetry by breaking the $\mathrm{SU}(2)$ symmetry of our quantum dynamics to $\mathrm{U}(1)$.
1 Introduction
@@
The growth of entanglement entropy also helps the understanding of open quantum many-body systems, which are ubiquitous in practical experiments due to the inevitable couplings to environments. Like the MPS representation for pure states, the density matrix of 1D open quantum many-body systems can be described by the matrix product density operators (MPDOs) [29, 30, 31]. The bipartition of MPDO through Schmidt decomposition further defines the entanglement entropy in operator space, which is known as the operator entanglement and characterizes the cost of an MPDO representation, i.e., how many Schmidt values are needed at least for faithfully representing an operator [32, 33, 34, 35, 36, 37], hence providing a measure for the classical simulability of open quantum many-body systems via the MPDO tensor-network method.
@@
In this work, we numerically study the far-from-equilibrium dynamics of operator entanglement in a dissipative quantum many-body system with $\mathrm{SU}(2)$ symmetry. We show that after the initial rise and fall, the operator entanglement also increases again in a logarithmic manner at long times in the $\mathrm{SU}(2)$-symmetric case. We find that this behavior can be fully understood from the corresponding $\mathrm{U}(1)$ subsymmetry by considering the symmetry-resolved operator entanglement [44, 45, 46]. But unlike the $\mathrm{U}(1)$-symmetric case, both the classical Shannon entropy and the symmetry-resolved operator entanglement now have nontrivial contributions to the late time logarithmic growth of operator entanglement. Our results show evidence that the logarithmic growth of operator entanglement at long times is a generic behavior of dissipative quantum many-body dynamics with $\mathrm{U}(1)$ as the symmetry or subsymmetry and for more broad dissipations beyond dephasing. We show the latter is even valid for open quantum systems with only U(1) symmetry by breaking the $\mathrm{SU}(2)$ symmetry of our quantum dynamics to $\mathrm{U}(1)$.
2 Model
@@
where the Hamiltonian $H=J\sum_{i}P_{i,i+1}$ with the exchange operator $P_{i,i+1}=2\mathbf{S}_{i}\cdot\mathbf{S}_{i+1}+1/2$ is the $1$D Heisenberg model with $\mathrm{SU}(2)$ symmetry. Here $\mathbf{S}_{i}$ denotes the spin-$1/2$ operator on site $i$, and $J$ is the nearest-neighbor spin coupling strength. We couple the system to an environment through the Lindblad operator $L_{i}=P_{i,i+1}$ with strength $\gamma$, which describes the dissipation proportional to the dipole interaction between neighbor sites; see Fig. 1(a) for the sketch. Similar model has been considered in Refs. [47, 48] to study the stability or absence of superdiffusion in a spin chain with fluctuating exchange couplings that break the integrability. Since $L_{i}$ commutes with the total spin operator, the above Lindblad master equation preserves the $\mathrm{SU}(2)$ symmetry. To study the quantum dynamics with $\mathrm{SU}(2)$ symmetry, we consider the initial state $\rho_{0}=|\psi_{0}\rangle\langle\psi_{0}|$ with $|\psi_{0}\rangle=\bigotimes_{i}(\left|\uparrow\right\rangle_{2i-1}\left|\downarrow\right\rangle_{2i}-\left|\downarrow\right\rangle_{2i-1}\left|\uparrow\right\rangle_{2i})/\sqrt{2}$, which is the product state of singlet pairs. This state has total spin 0 and is $\mathrm{SU}(2)$-symmetric.
Figure 1: Operator entanglement dynamics in the $\mathrm{SU}(2)$-symmetric dissipative quantum many-body system. (a) Sketch of the model. We consider a quantum spin chain with coherent nearest-neighbor coupling $J$ (blue arrows) and local dissipation proportional to the dipole interaction between neighbor sites with strength $\gamma$ (green arrows). (b) MPDO decomposition of the density matrix $\rho$ in terms of local tensors $\Gamma$ (blue squares) and $\lambda$ (orange circles). Here $\alpha_{i-1}$ and $\alpha_{i}$ are the bond indices, while $s_{i}$ denotes the combined physical index of the bra and ket legs at site $i$. The operator entanglement $S_{\mathrm{op}}$ at certain bond can be calculated from the Schmidt vectors $\lambda$. (c) Time evolution of $S_{\mathrm{op}}$ for the product initial state of singlet pairs with different dissipation strength $\gamma=0.05J$, $0.10J$, $0.15J$, $0.20J$, and $0.25J$. The black dashed line indicates the logarithmic growth of operator entanglement at long times (log-scale time axis), i.e., $S_{\mathrm{op}}(t\to\infty)=\eta\log_{2}(tJ)+S_{0}$. (d) Numerical prefactor $\eta$ and offset $S_{0}$ obtained from the local tangent of operator entanglement at time $t_{0}J$. The results are converged for time step $\delta tJ=0.5$ and maximum bond dimension $\chi=50000$.
3 MPDO decomposition of density matrix
@@
For the Liouvillian superoperator (1) that can be decomposed into terms involving at most two contiguous sites, the time evolution can be simulated using the infinite time-evolving block decimation (iTEBD) algorithm with a fourth-order Trotter decomposition of the matrix exponential of the Liouvillian $\exp(\mathcal{L}_{\sharp}\delta t)$ for a time step $\delta t$ [21]. We note that since the open quantum many-body dynamics governed by the Lindblad master equation is nonunitary, the canonical form of the infinite MPDO is in general not preserved in the original iTEBD algorithm [49]. As a consequence, the truncation errors accumulate fast and ruin the simulation at long times. To overcome this issue, we employ the improved algorithm for nonunitary evolutions proposed in Ref. [50], in which the tensors updated in each time step are reorthonormalized into the canonical form. To avoid the exponential growth of bond dimension, we also truncate the tensors $\{\Gamma^{[i]}\}$ and $\{\lambda^{[i]}\}$ at a maximum bond dimension $\chi$. The results shown in this work are all numerically converged.
4 Operator entanglement
@@
It is useful to also introduce the symmetry-resolved operator entanglement, which has attracted extensive attention recently [56, 57, 58, 59, 60, 61]. Specifically, following in Ref. [38], for a generic global symmetry $Q$, the operator entanglement for the MPDO decomposition (2) at certain bond can be split into different symmetry sectors, similarly to that of the state entanglement [44, 45, 46]. The symmetry-resolved operator entanglement in a specific symmetry sector $q$ is defined through the renormalized Schmidt coefficients $\hat{\lambda}_{q,i_{q}}=\lambda_{q,i_{q}}/\sqrt{p_{q}}$ belonging to that sector via $S_{\mathrm{op},q}\equiv-\sum_{i_{q}}\hat{\lambda}^{2}_{q,i_{q}}\log_{2}\hat{\lambda}^{2}_{q,i_{q}}$, where $i_{q}$ labels different Schmidt coefficients in the symmetry sector $q$ and $p_{q}=\sum_{i_{q}}\lambda^{2}_{q,i_{q}}$ is the probability of having symmetry charge $q$ in the half system such that $\sum_{i_{q}}\hat{\lambda}^{2}_{q,i_{q}}=1$. In our $\mathrm{SU}(2)$-symmetric case, different symmetry sectors are labeled by the half-system total spin $S$, and the probability of having total spin $S$ in the half system is $p_{S}=(2S+1)\sum_{i_{S}}\lambda_{(S,i_{S})}^{2}$ [note that in our notation for the $\mathrm{SU}(2)$ symmetry, each bond label $(S,i_{S})$ actually corespondents to $(2S+1)$ Schmidt coefficients with the same value $\lambda_{S,i_{S}}$]. Then the symmetry-resolved operator entanglement in the spin sector $S$ is given by $S_{\mathrm{op},S}=-(2S+1)\sum_{i_{S}}\hat{\lambda}_{(S,i_{S})}^{2}\log_{2}\hat{\lambda}_{(S,i_{S})}^{2}$ with $\hat{\lambda}_{(S,i_{S})}\equiv\lambda_{(S,i_{S})}/\sqrt{p_{S}}$. With this, the full operator entanglement can be recast as
@@
Figure 2: Symmetry-resolved operator entanglement in spin sectors. (a) Probabilities $p_{S}$ of the infinite chain with total spin $S$ in the half system at increasingly late times ($50\leq tJ\leq 250$ from light to dark) for $\gamma=0.05J$. The lines are fits with the trial function shown in the main text. (b) The corresponding symmetry-resolved operator entanglement $S_{\mathrm{op},S}$ as a function of time. Only the data with probabilities $p_{S}>10^{-4}$ are presented. (c) (d) The classical Shannon entropy $-\sum p_{S}\log_{2}p_{S}$ and the averaged symmetry-resolved operator entanglement $\sum p_{S}S_{\mathrm{op},S}$ as a function of time for $\gamma/J=0.05$, $0.10$, $0.15$, $0.20$, and $0.25$. The black dashed lines indicate the logarithmic growth at late times (log-scale time axis). The inserts present the corresponding prefactors. The results are converged for the time step $\delta tJ=0.5$ and maximum bond dimension $\chi=50000$.
5 Results
5.1 Logarithmic growth of operator entanglement
@@
5.2 Symmetry-resolved operator entanglement in spin sectors
@@
We also present the symmetry-resolved operator entanglement in different spin sectors in Fig. 2(b). Unlike Ref. [38], where the symmetry-resolved operator entanglement drops to very small values quickly, here the operator entanglement $S_{\mathrm{op},S}$ remains finite and large at late times, indicating that in addition to the classical Shannon entropy of probability distributions, the symmetry-resolved operator entanglement also have nontrivial contributions to the logarithmic growth of the total operator entanglement in our $\mathrm{SU}(2)$-symmetric case, as we show in Figs. 2(c) and 2(d). The nonvanishing contribution of symmetry-resolved operator entanglement at late times is compatible with the results shown in Refs. [42, 43], in which the authors studied the bipartite entanglement of the stationary states for unital quantum channels on a finite chain at $t\to\infty$ and found that the half-system operator entanglement scales logarithmically with the system size for the open quantum many-body systems with $\mathrm{SU}(2)$ symmetry. In this $t\to\infty$ limit, the nontrivial Hilbert space structure with dimension $12S+1>$ 1 for spin sectors $S>0$ leads to the divergent contributions from symmetry-resolved operator entanglement as the system size goes to infinity, otherwise the operator entanglement is fully captured by the classical correlations like the $\mathrm{U}(1)$-symmetric case considered in Ref. [38]. We note that although the operator entanglement shows a logarithmic behavior both in system size in the $t\to\infty$ limit and in time in the limit of infinite system size, it seems that these two phenomena do not have too much connection with each other as the prefactor $\eta$ in our work is different from those shown in Refs. [42, 43].
@@
5.3 Symmetry-resolved operator entanglement in magnetization sectors

### v3

@@ Abstract
The presence of symmetries can lead to nontrivial dynamics of operator entanglement in open quantum many-body systems, which characterizes the cost of an matrix product density operator (MPDO) representation of the density matrix in the tensor-network methods and provides insights into the corresponding classical simulability. One example is the $\mathrm{U}(1)$-symmetric open quantum systems with dephasing, in which the operator entanglement increases logarithmically at late times instead of being suppressed by the dephasing. Here we numerically study the far-from-equilibrium dynamics of operator entanglement in a dissipative quantum many-body system with the more complicated $\mathrm{SU}(2)$ symmetry and dissipations beyond dephasing. We show that after the initial rise and fall, the operator entanglement also increases again in a logarithmic manner at late times in the $\mathrm{SU}(2)$-symmetric case. We find that this behavior can be fully understood from the corresponding $\mathrm{U}(1)$ subsymmetry by considering the symmetry-resolved operator entanglement. Especially, the probability distribution of different $\mathrm{U}(1)$ sectors also follows the Gaussian distribution observed in the $\mathrm{U}(1)$-symmetric case with dephasing, with the variance growing in a power law. But unlike the latter, both the classical Shannon entropy associated with the probabilities for the half system being in different symmetry sectors and the corresponding symmetry-resolved operator entanglement now have nontrivial contributions to the late time logarithmic growth of operator entanglement in our $\mathrm{SU}(2)$-symmetric case. Our results show evidence that the logarithmic growth of operator entanglement at long times is a generic behavior of dissipative quantum many-body dynamics with $\mathrm{U}(1)$ as the symmetry or subsymmetry and for more broad dissipations beyond dephasing, although more analytical or numerical proofs are still required for this conjecture in the future studies. By breaking the $\mathrm{SU}(2)$ symmetry of our quantum many-body dynamics to $\mathrm{U}(1)$, we also show that the latter property—the logarithmic growth behavior of operator entanglement for more broad dissipations beyond dephasing—is valid even for open quantum systems with only $\mathrm{U}(1)$ symmetry.
I Introduction
@@
The growth of entanglement entropy also helps the understanding of open quantum many-body systems, which are ubiquitous in practical experiments due to the inevitable couplings to environments. Like the MPS representation for pure states, the density matrix of 1D open quantum many-body systems can be described by the matrix product density operators (MPDOs) [29, 30, 31]. The bipartition of MPDO through Schmidt decomposition further defines the entanglement entropy in operator space, which is known as the operator entanglement and characterizes the cost of an MPDO representation, i.e., how many Schmidt values are needed at least for faithfully representing an operator [32, 33, 34, 35, 36, 37], hence providing insights into the classical simulability of open quantum many-body systems via the MPDO tensor-network method.
@@
In this work, we numerically study the far-from-equilibrium dynamics of operator entanglement in a dissipative quantum many-body system with $\mathrm{SU}(2)$ symmetry. We show that after the initial rise and fall, the operator entanglement also increases again in a logarithmic manner at long times in the $\mathrm{SU}(2)$-symmetric case. We find that this behavior can be fully understood from the corresponding $\mathrm{U}(1)$ subsymmetry by considering the symmetry-resolved operator entanglement [44, 45, 46]. Especially, the probability distribution of different $\mathrm{U}(1)$ sectors also follows the Gaussian distribution observed in the $\mathrm{U}(1)$-symmetric case with dephasing [38], with the variance growing in a power law. But unlike the latter, both the classical Shannon entropy and the symmetry-resolved operator entanglement now have nontrivial contributions to the late-time logarithmic growth of operator entanglement in our $\mathrm{SU}(2)$-symmetric dissipative quantum many-body dynamics. Our results show evidence that the logarithmic growth of operator entanglement at long times is a generic behavior of dissipative quantum many-body dynamics with $\mathrm{U}(1)$ as the symmetry or subsymmetry and for more broad dissipations beyond dephasing, although more analytical or numerical proofs are still required for this conjecture in the future studies. By breaking the $\mathrm{SU}(2)$ symmetry of our quantum many-body dynamics to $\mathrm{U}(1)$, we also show that the latter property—the logarithmic growth behavior of operator entanglement for more broad dissipations beyond dephasing—is valid even for open quantum systems with only $\mathrm{U}(1)$ symmetry.
The remaining part of this article is organized as follows. In Sec. II, we introduce the dissipative quantum many-body model studied in this work. Then in Sec. III, we describe the MPDO decomposition of density matrix. After introducing the (symmetry-resolved) operator entanglement in Sec. IV, we present the numerical results in Sec. V. Finally, the conclusion is provided in Sec. VI.
II Model
@@
where the Hamiltonian $H=J\sum_{i}P_{i,i+1}$ with the exchange operator $P_{i,i+1}=2\mathbf{S}_{i}\cdot\mathbf{S}_{i+1}+1/2$ is the 11D Heisenberg model with $\mathrm{SU}(2)$ symmetry. Here $\mathbf{S}_{i}$ denotes the spin-$1/2$ operator on site $i$, and $J$ is the nearest-neighbor spin coupling strength. We couple the system to an environment through the Lindblad operator $L_{i}=P_{i,i+1}$ with strength $\gamma$, which describes the dissipation proportional to the dipole interaction between neighbor sites; see Fig. 1(a) for the sketch. Similar model has been considered in Refs. [47, 48] to study the stability or absence of superdiffusion in a spin chain with fluctuating exchange couplings that break the integrability. Since $L_{i}$ commutes with the total spin operator, the above Lindblad master equation preserves the $\mathrm{SU}(2)$ symmetry. To study the quantum dynamics with $\mathrm{SU}(2)$ symmetry, we consider the initial state $\rho_{0}=|\psi_{0}\rangle\langle\psi_{0}|$ with $|\psi_{0}\rangle=\bigotimes_{i}(\left|\uparrow\right\rangle_{2i-1}\left|\downarrow\right\rangle_{2i}-\left|\downarrow\right\rangle_{2i-1}\left|\uparrow\right\rangle_{2i})/\sqrt{2}$, which is the product state of singlet pairs. This state has total spin 0 and is $\mathrm{SU}(2)$-symmetric.
Figure 1: Operator entanglement dynamics in the $\mathrm{SU}(2)$-symmetric dissipative quantum many-body system. (a) Sketch of the model. We consider a quantum spin chain with coherent nearest-neighbor coupling $J$ (blue arrows) and local dissipation proportional to the dipole interaction between neighbor sites with strength $\gamma$ (green arrows). (b) MPDO decomposition of the density matrix $\rho$ in terms of local tensors $\Gamma$ (blue squares) and $\lambda$ (orange circles). Here $\alpha_{i-1}$ and $\alpha_{i}$ are the bond indices, while $s_{i}$ denotes the combined physical index of the bra and ket legs at site ii. The operator entanglement $S_{\mathrm{op}}$ at certain bond can be calculated from the Schmidt vectors $\lambda$. (c) Time evolution of $S_{\mathrm{op}}$ for the product initial state of singlet pairs with different dissipation strength $\gamma=0.05J$, $0.10J$, $0.15J$, $0.20J$, and $0.25J$. The black dashed line indicates the logarithmic growth of operator entanglement at long times (log-scale time axis), i.e., $S_{\mathrm{op}}(t\to\infty)=\eta\log_{2}(tJ)+S_{0}$. (d) Numerical prefactor $\eta$ and offset $S_{0}$ obtained from the local tangent of operator entanglement at time $t_{0}J$. The results are converged for time step $\delta tJ=0.5$ and maximum bond dimension $\chi=50000$.
III MPDO decomposition of density matrix
@@
For the Liouvillian superoperator (1) that can be decomposed into terms involving at most two contiguous sites, the time evolution can be simulated using the infinite time-evolving block decimation (iTEBD) algorithm with a fourth-order Trotter decomposition of the matrix exponential of the Liouvillian $\exp(\mathcal{L}_{\sharp}\delta t)$ for a time step $\delta t$ [21]. We note that since the open quantum many-body dynamics governed by the Lindblad master equation is nonunitary, the canonical form of the infinite MPDO is in general not preserved in the original iTEBD algorithm [49]. As a consequence, the truncation errors accumulate fast and ruin the simulation at long times. To overcome this issue, we employ the improved algorithm for nonunitary evolutions proposed in Ref. [50], in which the tensors updated in each time step are reorthonormalized into the canonical form. To avoid the exponential growth of bond dimension, we also truncate the tensors $\{\Gamma^{[i]}\}$ and $\{\lambda^{[i]}\}$ at a maximum bond dimension $\chi$. The results shown in this work are all numerically converged; see Appendix A.
IV Operator entanglement
@@
It is useful to also introduce the symmetry-resolved operator entanglement, which has attracted extensive attention recently [38, 44, 45, 46]. In this work, we follow the definition proposed in Refs. [38, 44], while alternative definitions are also possible [45, 46]. Specifically, for a generic global symmetry $Q$, the operator entanglement for the MPDO decomposition (2) at certain bond can be split into different symmetry sectors, similar to that of the state entanglement [56, 57, 58, 59, 60, 61]. The symmetry-resolved operator entanglement in a specific symmetry sector $q$ is defined through the renormalized Schmidt coefficients $\hat{\lambda}_{q,i_{q}}=\lambda_{q,i_{q}}/\sqrt{p_{q}}$ belonging to that sector via $S_{\mathrm{op},q}\equiv-\sum_{i_{q}}\hat{\lambda}^{2}_{q,i_{q}}\log_{2}\hat{\lambda}^{2}_{q,i_{q}}$, where $i_{q}$ labels different Schmidt coefficients in the symmetry sector $q$ and $p_{q}=\sum_{i_{q}}\lambda^{2}_{q,i_{q}}$ is the probability of having symmetry charge $q$ in the half system such that $\sum_{i_{q}}\hat{\lambda}^{2}_{q,i_{q}}=1$. In our $\mathrm{SU}(2)$-symmetric case, different symmetry sectors are labeled by the half-system total spin $S$, and the probability of having total spin $S$ in the half system is $p_{S}=(2S+1)\sum_{i_{S}}\lambda_{(S,i_{S})}^{2}$ [note that in our notation for the $\mathrm{SU}(2)$ symmetry, each bond label $(S,i_{S})$ actually corespondents to $(2S+1)$ Schmidt coefficients with the same value $\lambda_{S,i_{S}}$]. Then the symmetry-resolved operator entanglement in the spin sector $S$ is given by $S_{\mathrm{op},S}=-(2S+1)\sum_{i_{S}}\hat{\lambda}_{(S,i_{S})}^{2}\log_{2}\hat{\lambda}_{(S,i_{S})}^{2}$ with $\hat{\lambda}_{(S,i_{S})}\equiv\lambda_{(S,i_{S})}/\sqrt{p_{S}}$. With this, the full operator entanglement can be recast as
@@
Figure 2: Symmetry-resolved operator entanglement in spin sectors. (a) Probabilities $p_{S}$ of the infinite chain with total spin $S$ in the half system at increasingly late times ($50\leq tJ\leq 250$ from light to dark) for $\gamma=0.05J$. The lines are fits with the trial function shown in the main text. (b) The corresponding symmetry-resolved operator entanglement $S_{\mathrm{op},S}$ as a function of time. Only the data with probabilities $p_{S}>10^{-4}$ are presented. (c) (d) The classical Shannon entropy $\sum p_{S}\log_{2}p_{S}$ and the averaged symmetry-resolved operator entanglement $\sum p_{S}S_{\mathrm{op},S}$ as a function of time for $\gamma/J=0.05$, $0.10$, $0.15$, $0.20$, and $0.25$. The black dashed lines indicate the logarithmic growth at late times (log-scale time axis). The inserts present the corresponding prefactors. The results are converged for the time step $\delta tJ=0.5$ and maximum bond dimension $\chi=50000$.
V Results
V.1 Logarithmic growth of operator entanglement
@@
V.2 Symmetry-resolved operator entanglement in spin sectors
@@
We also present the symmetry-resolved operator entanglement in different spin sectors in Fig. 2(b). Unlike Ref. [38], where the symmetry-resolved operator entanglement drops to very small values quickly, here the operator entanglement $S_{\mathrm{op},S}$ remains finite and large at late times, indicating that in addition to the classical Shannon entropy of probability distributions, the symmetry-resolved operator entanglement also have nontrivial contributions to the logarithmic growth of the total operator entanglement in our $\mathrm{SU}(2)$-symmetric case, as we show in Figs. 2(c) and 2(d). The nonvanishing contribution of symmetry-resolved operator entanglement at late times is compatible with the results shown in Refs. [42, 43], in which the authors studied the bipartite entanglement of the stationary states for unital quantum channels on a finite chain at $t\to\infty$ and found that the half-system operator entanglement scales logarithmically with the system size for the open quantum many-body systems with $\mathrm{SU}(2)$ symmetry. In this $t\to\infty$ limit, it was shown that the total operator entanglement involves both quantum and classical contributions; see Eq. (23) in Ref. [43] or Eq. (20) in Ref. [42]. The former is related to the nontrivial Hilbert space structure with dimension $2S+1>1$ for each spin sectors $S>0$, while the latter is attributed to the probability distribution of states among different spin sectors. Both the quantum and classical contributions were shown to scale logarithmically with the system size and diverge in the thermodynamic limit, which is compatible with our results that both the classical Shannon entropy of probability distributions and the averaged symmetry-resolved operator entanglement for an infinite spin chain grow logarithmically with the time and also diverge as $t\to\infty$. Therefore, it would be expected that the nonvanishing contribution of symmetry-resolved operator entanglement at late times in our $\mathrm{SU}(2)$-symmetric case arises from the nontrivial spin sector structures. We note that although the operator entanglement shows a logarithmic behavior both in system size in the $t\to\infty$ limit and in time in the limit of infinite system size, it seems that these two phenomena do not have too much connection with each other as the prefactor $\eta$ in our work is different from those shown in Refs. [42, 43].
@@
V.3 Symmetry-resolved operator entanglement in magnetization sectors